The current in a single-loop circuit containing a single resistance R is 6 amps. When an additional resistance of 12 Ω is inserted in series with R, the current drops to 3 amps. What is R?
Answer #1

Solution)

Given,

Current, I1=6 A

Resistance, R2=12 ohms

Current, I2=3 A

According to Ohm's law, we can conclude

I1*R1= I2*R2

Substitute values,

6*R=3*(R+12)

Since, 12 ohms connected in series with R

So, 6R=3R + 36

3R=36

Resistance, R=12 ohms (Ans)

==========